# USP-Gaussian: Unifying Spike-based Image Reconstruction, Pose Correction and Gaussian Splatting

### v1

@@ 5.1 Datasets
Synthetic Dataset. We construct the synthetic dataset based on scenes from Deblur-NeRF [15] for quantitative evaluation. Following the motion trajectory specified in Deblur-NeRF, we start with rendering sharp images at a resolution of $400\times 600$ for each camera viewpoint, with $12$ additional intermediate frames inserted between consecutive views. Subsequently, we employ the XVFI [22] to insert $7$ images between each frame pair. We employ the spike simulator on the enhanced video sequence to emulate the real-world spike stream as in prior research [3], thus generating a spike stream of $97$ frames for each viewpoint.
Real-world Dataset. To construct the real-world dataset, we shake the spike camera extremely fast while maintaining a fixed forward-facing orientation and extract $97$ spike frames per view, consistent with the configuration in the synthetic dataset. For pose estimation, we convert the raw spike stream in each view to the short-exposure image as defined in Eq. 6 and subsequently employ COLMAP [21] to estimate the pose sequence. Nevertheless, the resulting pose estimates are notably inaccurate due to the substantial noise encoded in the reconstructed sequence.
@@ 5.2 Experimental Results
We compare our USP-Gaussian with two-stage cascading pipelines and approaches specially designed for the spike camera, i.e., SpikeNeRF [45] and SpikeGS [32]. For the two-stage approach, we initially apply TFP [43], TFI [43], and Spk2ImgNet [36] to reconstruct sharp images from the input spike, followed by the common 3DGS processing pipeline. All Gaussian-based methods are conducted based on the same 3DGS backbone, and identical camera poses input for a fair comparison. While SpikeGS requires several poses per spike stream for the exposure loss, which presents a notable challenge for COLMAP in real-world data, we omit the exposure loss and retain the rest of the framework. All comparative metrics are calculated based on the 3D reconstruction task, with further comparison on the Spike-to-Image reconstruction task and implementation details available in the supplementary materials.
Synthetic Dataset. Quantitative and qualitative comparisons on the synthetic dataset are shown in Tab. 1 and Fig. 4 respectively. As illustrated in the Tab. 1, our USP-Gaussian achieves the best performance across most metrics in various scenes. Two-stage cascading methods TFP-3DGS and TFI-3DGS yield poor restoration due to the substantial noise embedded in the initially reconstructed sequence. Similarly, although Spk2ImgNet performs well on its original REDS dataset [36], it experiences substantial degradation due to the dataset domain gap. SpikeGS, as the primary comparison algorithm in this paper, heavily relies on the BSN [13] for self-supervised Spike-to-Image reconstruction. Nevertheless, the denoising performance of the BSN deteriorates in the presence of substantial noise, leading to pronounced grid artifacts and the attenuation of high-frequency details of the SpikeGS. SpikeNeRF faces a comparable challenge, as it requires noisy spike images as input and relies on SNN and NeRF for denoising.
@@
Real-world Dataset. While we have demonstrated the advantages of joint learning between 3DGS and Recon-Net on the synthetic dataset, we further conduct the visual comparison on real-world data in Fig. 5, to highlight the necessity of correcting pose during the optimization. Due to the rapid camera ego-motion and the low imaging quality of the spike camera characterized by low spatial resolution, insufficient texture details, and underexposure, the pose sequence estimated by the COLMAP is highly inaccurate, presenting significant challenges for sequential-based methods to recover captured 3D scene. As observed in Fig. 5, baseline methods fail to reconstruct the corresponding scene owing to the improper pose representation, while our USP-Gaussian overcomes this limitation by benefiting from the unifying optimization of the pose.
@@
$\mathbb{L}_{\text{recon}}$ $\mathcal{L}_{\text{recon}}$ PSNR SSIM LPIPS PSNR SSIM LPIPS
@@ 5.3 Ablation Study
Joint Learning. The highlight of USP-Gaussain lies in the design of the joint loss, which aligns outputs from the 3DGS and Recon-Net thus achieving complementary performance improvement. We conduct the ablation experiment on loss functions of $\mathcal{L}_{\text{gs}}$, $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{joint}}$ with quantitative results listed in Tab. 2. As illustrated in the table, the incorporation of the joint loss (ID-V) yields notable improvements in PSNR (18.24%/11.89%), SSIM (14.93%/20.76%), and LPIPS (15.38%/50.28%) for both 3DGS and Recon-Net compared to independent optimization ((ID-I and ID-III).
Further visual ablation comparisons of the restored image and the training curve are presented in Fig. 1. Visual comparison substantiates that the joint optimization of GS and Recon-Net significantly enhances the texture details for both two. Regarding the training curve, the independent 3DGS training (ID-I) demonstrates slow convergence due to the absence of crucial information in the long-exposure images. Concurrently, the independent training of Recon-Net (ID-III) tends to overfit due to the simplicity of the spike reconstruction task coupled with the limited constraint from self-supervised loss. To this end, our proposed joint optimization framework provides synergistic benefits for both, i.e., Recon-Net effectively compensates for deficient motion features embedded in the long-exposures image while the multi-view consistency afforded by 3DGS culminates in coherent outputs from the Recon-Net. The training curve in Fig. 1 further demonstrates the synergistic optimization process of Recon-Net and 3DGS.
@@
Multi-reblur Loss. The multi-reblur loss $\mathcal{L}_{\text{recon}}$ in Eq. 7 is designed to substitute the single-reblur loss $\mathbb{L}_{\text{recon}}$ in Eq. 8. Without further constraints, $\mathbb{L}_{\text{recon}}$ will lead Recon-Net to learn an identity mapping from the input spike stream to the corresponding long-exposure image as demonstrated in [3]. Comparative analysis in Tab. 2 depicts that $\mathcal{L}_{\text{recon}}$ (ID-V) yields beneficial improvement than $\mathbb{L}_{\text{recon}}$(ID-IV). Further comparison on the training curve in Fig. 6 reveals the mechanism of collaborative training, i.e., the multi-reblur loss enhances the performance of Recon-Net, thereby facilitating the 3DGS training.
Complementary Input. We utilize complementary long and short spike stream inputs for Recon-Net to address the fixed short window limitation inherent in previous methods [36, 31, 4]. As illustrated in the visual comparison in Fig. 7, within the same joint optimization framework, the inclusion of long spike stream inputs effectively suppresses salt-and-pepper noise in the output of the reconstruction network, leading to improved SSIM and LPIPS metrics.

### v2

@@ 5.1 Datasets
Synthetic Dataset. We construct the synthetic dataset based on Deblur-NeRF [19] scenes for quantitative evaluation. We start with rendering $13$ sharp images for each viewpoint, with further interpolation algorithm XVFI [25] to insert $7$ images between each frame pair. We apply the spike simulator on the interpolated sequence to emulate the real-world spike as in [3], thus generating $97$ spike frames for each viewpoint.
Real-world Dataset. To construct the real-world dataset, we shake the spike camera extremely fast while maintaining a fixed forward-facing orientation and extract $97$ spike frames per view, consistent with the configuration in the synthetic dataset. For pose estimation, we convert the raw spike stream in each view to the short-exposure image as defined in Eq. 6 and subsequently employ COLMAP [24] to estimate the pose sequence.
@@ 5.2 Experimental Results
We compare our USP-Gaussian with two-stage cascading pipelines and approaches specially designed for the spike camera, i.e., SpikeNeRF [50] and SpikeGS [37]. For the two-stage approach, we initially apply TFP [48], TFI [48], and Spk2ImgNet [41] to reconstruct sharp images from the input spike followed by the common 3DGS processing pipeline, with the code implemented by Spike-Zoo [4]. All Gaussian-based methods are conducted based on the same 3DGS backbone, and identical camera poses input for a fair comparison. While SpikeGS requires several poses per spike stream for the exposure loss, which presents a notable challenge for COLMAP in real-world data, we omit the exposure loss and retain the rest of the framework. All comparative metrics are calculated based on the 3D reconstruction task, with further comparison on the spike-to-image reconstruction task and implementation details available in the supplementary materials.
Table 2: Comparison of our method with SpikeGS on the wine dataset with inaccurate perturbed poses.
Metrics SpikeGS Ours
0% 10% 20% 30% 0% 10% 20% 30%
PSNR $\uparrow$ 25.69 18.83 17.75 16.44 26.47 26.05 25.93 23.46
SSIM $\uparrow$ 0.835 0.607 0.592 0.567 0.862 0.854 0.851 0.772
LPIPS $\downarrow$ 0.247 0.372 0.411 0.446 0.176 0.184 0.186 0.245
Synthetic Dataset. Quantitative and qualitative comparisons on the synthetic dataset are shown in Tab. 1 and Fig. 4 respectively. As illustrated in the Tab. 1, our USP-Gaussian achieves the best performance across most metrics in various scenes. Two-stage cascading methods TFP-3DGS and TFI-3DGS yield poor restoration due to the substantial noise embedded in the initially reconstructed sequence. Similarly, although Spk2ImgNet performs well on its original REDS dataset [41], it experiences substantial degradation due to the dataset domain gap. SpikeGS, as the primary comparison algorithm in this paper, heavily relies on the BSN [15] for self-supervised Spike-to-Image reconstruction. Nevertheless, the denoising performance of the BSN deteriorates in the presence of substantial noise, leading to pronounced grid artifacts and the attenuation of high-frequency details of the SpikeGS. SpikeNeRF faces a comparable challenge, as it requires noisy spike images as input and relies on the SNN for denoising.
@@
In the previous part, accurate poses are obtained by applying COLMAP on sharp images rendered by Blender. To demonstrate that the integration of pose correction enhances the robustness of our method, we apply random perturbations of 10%, 20%, and 30% to the accurate poses in the wine scene, generating inaccurate poses for SpikeGS and our method with the comparison shown in Tab. 2.
Figure 6: Visual comparison of initial and optimized poses.
Table 3: Translation and rotation poses error comparison.
State Translation $\downarrow$ Rotation $\downarrow$
10% 20% 30% Avg 10% 20% 30% Avg
Initial 1.302 2.967 4.193 2.820 1.132 2.150 3.159 2.147
Optimized 0.855 1.781 1.493 1.376 0.174 0.320 1.285 0.593
$\Delta$ (%) 34.3% 39.9% 64.4% 51.3% 84.6% 85.1% 59.2% 72.4%
Furthermore, visual trajectories and quantitative errors of the initial noisy, and optimized poses are depicted in Fig. 6 and Tab. 3 respectively, with MAE measuring translation error and radians measuring rotation error, demonstrating the pose correction capability of our USP-Gaussian.
Real-world Dataset. We also conduct the visual comparison on the real-world dataset as shown in Fig. 5. Due to the rapid camera ego-motion and the low imaging quality of the spike camera characterized by low spatial resolution, insufficient texture details, and underexposure, the pose sequence estimated by the COLMAP is highly inaccurate, presenting significant challenges for previous methods to recover accurate 3D representation as observed in Fig. 5. Nevertheless, our USP-Gaussian overcomes this limitation benefiting from the unifying optimization of the pose.
@@
$\mathbb{L}_{\text{rec}}$ $\mathcal{L}_{\text{rec}}$ PSNR SSIM LPIPS PSNR SSIM LPIPS
@@ 5.3 Ablation Study
I. Joint Learning. The highlight of USP-Gaussain lies in the design of the joint loss, which aligns outputs from the 3DGS and Recon-Net thus achieving complementary performance improvement. We conduct the ablation experiment on loss functions of $\mathcal{L}_{\text{gs}}$, $\mathcal{L}_{\text{rec}}$ and $\mathcal{L}_{\text{joint}}$ with quantitative results listed in Tab. 4, which shows that the incorporation of the joint loss (ID-V) yields notable improvements for both 3DGS and Recon-Net compared to independent optimization (ID-I and ID-III).
Further visual ablation comparisons of the restored image and the training curve are presented in Fig. 1. Visual comparison substantiates that the joint optimization of 3DGS and Recon-Net significantly enhances the texture details for both two. Regarding the training curve, the independent 3DGS training (ID-I) demonstrates slow convergence due to the absence of crucial information in the long-exposure images. Concurrently, the independent training of Recon-Net (ID-III) tends to overfit due to the simplicity of the spike reconstruction task. To this end, our proposed joint optimization framework provides synergistic benefits for both, i.e., Recon-Net effectively compensates for deficient motion features embedded in the long-exposures image while the multi-view consistency afforded by 3DGS culminates in coherent outputs from the Recon-Net.
@@
II. Multi-reblur Loss. The multi-reblur loss $\mathcal{L}_{\text{rec}}$ in Eq. 7 is designed to substitute the single-reblur loss $\mathbb{L}_{\text{rec}}$ in Eq. 8. Without further constraints, $\mathbb{L}_{\text{rec}}$ will lead Recon-Net to learn an identity mapping from the input spike stream to the corresponding long-exposure image as demonstrated in [3]. Comparative analysis in Tab. 4 depicts that $\mathcal{L}_{\text{rec}}$ (ID-V) yields beneficial improvement than $\mathbb{L}_{\text{rec}}$(ID-IV). Further comparison on the training curve in Fig. 7 reveals the mechanism of collaborative training, i.e., the multi-reblur loss enhances the performance of Recon-Net, thereby facilitating the training of the 3DGS.
III. Complementary Input. We utilize complementary long and short spike stream inputs for Recon-Net to address the fixed short window limitation inherent in previous methods [41, 36, 6]. As illustrated in the visual comparison in Fig. 8, the inclusion of the long spike effectively suppresses salt-and-pepper noise in the Recon-Net output, leading to improved SSIM and LPIPS metrics.
@@ 6 Conclusion
Limitation. The limitation of USP-Gaussian is that it requires longer training time and more GPU memory compared to previous methods due to the additional poses and Recon-Net parameters to be optimized.
Acknowledgments
We sincerely appreciate Yuyan Chen (HUST) for her valuable suggestions and for polishing the figures. This work was supported by the National Natural Science Foundation of China (62422601, 62176003, 62088102, 62306015), the China Postdoctoral Science Foundation (2023T160015), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Beijing Municipal Science and Technology Program (Z241100004224004) and the Beijing Nova Program (20230484362,20240484703).